OpenStudy (anonymous):

Help. This one lost me. Solve the formula for the specified variable. I=nE/nr+R. Solve for n

OpenStudy (anonymous):

Is that a one on the left side of equation or an l if it is an l this is \[1=nE/(nr+R)\] multiply the denominator over nrl+Rl=nE get all the terms with n on one side nrl-nE=-Rl factor out n n(rl-E)=-Rl divide n=-Rl/(rl-E)

OpenStudy (anonymous):

HI, the only possible answers are: A. n = IR(Ir - E) B. n = -R/Ir-E C. n = IR/Ir+E D. n = -IR/Ir-E

OpenStudy (anonymous):

I'm sorry I reversed the sign on accident in the denominator. The answer is D. Does what I id make sense?

OpenStudy (anonymous):

Yes, thanks for your help. I have a picture of a graph also that I am supposed to find the equation from. Do you know of any good graph calculators?

OpenStudy (anonymous):

I use a ti-89 but anything from ti-82 on would be good. You could also get computer software for graphing

OpenStudy (anonymous):

Can you take a look at this one: Evaluate the function for the given value of x. f(x)=-6x^-4x-4, x=-2 I came up with f(-2)=-20. Does that look right?

OpenStudy (anonymous):

You get \[-6(-2)^4-4(-2)-4\] which simplifies to \[-6(16)+8-4\] so I would check your work again

OpenStudy (anonymous):

Possible answers are: A) f(-2)=-36. B) f(-2)=-17 C) f(-2)=-20 D) f(-2)=-28

OpenStudy (anonymous):

I'm sorry, maybe I misread your equation. Could you repost it please.

OpenStudy (anonymous):

Evaluate the function for the given value of x. f(x)=-6x^-4x-4, x=-2 I came up with f(-2)=-20

OpenStudy (anonymous):

I'm assuming that the ^ symbol means squared on the keyboard.

OpenStudy (anonymous):

not necessarily squared but "to the power"

OpenStudy (anonymous):

How do I make the little 2 above and to the right of the x

OpenStudy (anonymous):

That's what it's supposed to be in the expression

OpenStudy (anonymous):

^2

OpenStudy (anonymous):

but from there you get \[-6(4)+8-4=-24+8-4=-20\] so you are correct

OpenStudy (anonymous):

Ok, thanks, I did something right! If you don't mind looking at one more... Graph the exponential function and state the domain and range. y=5(x-3). I can't copy the graphs on here, but the possible answers to the domains/ranges are: A) Domain: (-8,8), Range: (-8,0) B) Domain: (0,8), Range: (0,8) C) Domain: (-8,0), Range: (-8,0) D) Domain: (-8,8) Range: (0,8)

OpenStudy (anonymous):

I'm not sure you copied the right equation there. The domain and range for that equation is all real numbers

OpenStudy (anonymous):

Lol, it's really on a workbook: Graph the exponential function and state the domain and range. y=5(x-3).

OpenStudy (anonymous):

And those I listed are the only options for an answer, besides the pics of the graphs. Wish I could figure out how to copy/paste those...

OpenStudy (anonymous):

is it \[5^{x-3}\]?

OpenStudy (anonymous):

That's not the way it's written, it looks just exactly like I typed it, but maybe it's supposed to be the way you wrote it?? Your way, does it have a solution to one of those I listed?

OpenStudy (anonymous):

I picked B and it was wrong.

OpenStudy (anonymous):

I am really unsure, sorry

OpenStudy (anonymous):

That's ok, I really do appreciate your help. I have other's but still trying to solve on my own at the moment. I will get back on later if you happen to still be here. Thanks again.

OpenStudy (anonymous):

okay you can message me if need be

OpenStudy (anonymous):

I don't know how to message you. First time on this site. How do I reach you?